# iVideoGPT: Interactive VideoGPTs are Scalable World Models

### v2

@@ Abstract
World models empower model-based agents to interactively explore, reason, and plan within imagined environments for real-world decision-making. However, the high demand for interactivity poses challenges in harnessing recent advancements in video generative models for developing world models at scale. This work introduces Interactive VideoGPT (iVideoGPT), a scalable autoregressive transformer framework that integrates multimodal signals—visual observations, actions, and rewards—into a sequence of tokens, facilitating an interactive experience of agents via next-token prediction. iVideoGPT features a novel compressive tokenization technique that efficiently discretizes high-dimensional visual observations. Leveraging its scalable architecture, we are able to pre-train iVideoGPT on millions of human and robotic manipulation trajectories, establishing a versatile foundation that is adaptable to serve as interactive world models for a wide range of downstream tasks. These include action-conditioned video prediction, visual planning, and model-based reinforcement learning, where iVideoGPT achieves competitive performance compared with state-of-the-art methods. Our work advances the development of interactive general world models, bridging the gap between generative video models and practical model-based reinforcement learning applications.
@@ 1 Introduction
Despite the fundamental connection, significant gaps remain between the development of generative models for video generation and world models for agent learning. One primary challenge is achieving the best of both interactivity and scalability. In the realm of model-based RL, world models predominantly utilize recurrent network architecture. This design naturally allows the transition of observations or latent states conditioned on actions in each step, facilitating interactive behavior learning [28, 77, 32]. However, these models mostly focus on games or simulated environments with visually simple data and have limited capability to model complex, in-the-wild data at scale. In contrast, Internet-scale video generative models [35, 6, 10] can synthesize realistic long videos that are controllable via text descriptions [104] or future action sequences [97]. Although such models allow for high-level, long-horizon planning [18], their trajectory-level interactivity does not provide sufficient granularity for agents to efficiently learn precise behavior as basic skills step-by-step.
@@
In this work, we explore the development of world models that are both interactive and scalable within a GPT-like autoregressive transformer framework [86, 72]. Pioneering efforts have been made recently through diffusion models [98] and masked generative models [11]. Nevertheless, utilizing autoregressive transformers for world models offers distinct advantages such as seamless integration with the established Large Language Model (LLM) ecosystem [105] and greater flexibility in handling diverse conditions without the need for specific architectural modifications like adapter modules [74, 102]. We present Interactive VideoGPT (iVideoGPT), a scalable world model architecture that incorporates multimodal signals, including visual observations, actions, and rewards, in an interactively autoregressive manner. Unlike multimodal LLMs that discretize visual observations into tokens frame-by-frame using image tokenizers [53], a key innovation of iVideoGPT is to learn compressive tokenization for each observation conditioned on rich contextual observations, achieving an asymptotic 16$\times$ reduction in token sequence length. We highlight that more compact video tokenization could not only facilitate more efficient training and generation but also enhance video quality. This is achieved by decoupling context from dynamics, allowing the model to focus on predicting the motion of objects while maintaining temporal consistency within the scene [95].
@@
We demonstrate a series of practical applications of iVideoGPT for visual robotic manipulation, as illustrated in Figure 1. Mirroring the two-phase approach popularized by Large Language Models (LLMs), our methodology involves pre-training followed by domain-specific adaptation. During pre-training, iVideoGPT is scalable for action-free video prediction on a mixture of over one million trajectories from both robotic and human manipulation [67, 24]. The pre-trained iVideoGPT serves as a versatile foundation that can be adapted into interactive world models for various downstream tasks. These include action-conditioned video prediction [20, 15], visual planning [82], and visual model-based RL [101]. Furthermore, we showcase the pre-trained transformer’s preliminary capability in iVideoGPT for zero-shot video generation without fine-tuning, requiring only the tokenizer to be adapted to unseen domains.
@@
Figure 1: Practical applications of iVideoGPT, which is designed to scale, allowing pre-training on millions of human and robotic manipulation trajectories. This results in a versatile foundation of interactive world models, adaptable to a wide range of downstream tasks.
@@
  • We pre-train iVideoGPT on a large-scale dataset comprising millions of robotic and human manipulation trajectories and adapt it to domain-specific tasks. The pre-trained model weights will be made available to encourage further research.
@@ 2 Problem Formulation
As depicted in Figure 2, a majority of advanced video generation models [97, 7, 100], including VideoGPT, can not deal with the interactive video prediction problem because they design non-causal modules fusing information along the temporal dimension. Existing world models in the literature of MBRL [28, 77], such as Dreamer, utilize recurrent architecture but lack scalability.
@@ 3 Interactive VideoGPT
In this section, we introduce Interactive VideoGPT, a scalable world model architecture with great flexibility to integrate multimodal signals, including visual observations, actions, rewards, and other potential sensory inputs. At its core, iVideoGPT consists of a compressive tokenizer to discretize video frames and an autoregressive transformer predicting subsequent tokens (Section 3.1). This model can acquire broad world knowledge through pre-training on diverse video data (Section 3.2) and then effectively transfer to downstream tasks incorporating additional modalities (Section 3.3).
@@ 3.1 Architecture
After tokenization, the video is flattened into a sequence of tokens: $x=(z_{1}^{(1)},\dots,z_{1}^{(N)},\text{{[S]}},z_{2}^{(1)},\dots,z_{2}^{(N)},\dots,\text{{[S]}},z_{T_{0}+1}^{(1)},\dots,z_{T_{0}+1}^{(n)},\dots)$ with a length of $L=(N+1)T_{0}+(n+1)(T-T_{0})-1$. Special slot tokens [S] are inserted to delineate frame boundaries and facilitate the integration of extra low-dimensional modalities such as actions (see Section 3.3 for details). As Figure 3b, a GPT-like autoregressive transformer is utilized for interactive video prediction through next-token generation frame-by-frame. In this work, we take the model size of GPT-2 [73] but adopt the LLaMA architecture [83] in order to embrace the latest innovations for LLM architecture, such as rotary position embedding [81].
@@ 3.2 Pre-Training
While there are numerous videos available on the Internet, due to computational limitations, we specifically pre-train iVideoGPT for the robotic manipulation domain. We leverage a mixture of 35 datasets from the Open X-Embodiment (OXE) dataset [67] and the Something-Something v2 (SSv2) dataset [24], totaling 1.5 million trajectories (see Appendix A.2 for details). OXE is a diverse collection of robot learning datasets from a variety of robot embodiments, scenes, and tasks. These datasets are highly heterogeneous but can be easily unified in the action-free video prediction task. To further enhance the diversity, we also include SSv2, a dataset of human-object interaction videos, as previous work has demonstrated knowledge transfer from these human manipulation videos for learning a world model for robotic manipulation tasks [95, 59].
@@ 3.3 Fine-Tuning
Our architecture is designed to flexibly incorporate additional modalities for learning interactive world models, as illustrated in Figure 3b. Actions are integrated by linear projection and adding to the slot token embeddings. For reward prediction, instead of learning independent reward predictors, we add a linear head to the last token’s hidden state of each observation. This multi-task learning approach can enhance the model’s focus on task-relevant information, thereby improving prediction accuracy for control tasks [55]. We use a mean-squared error loss for reward prediction in addition to the cross-entropy loss in Eq. (3).
@@
We choose to update the full model, including the tokenizer, for downstream tasks, finding this strategy more effective than parameter-efficient fine-tuning methods [37]. This is likely due to the limited diversity of our pre-trained data compared to Internet-scale images, which, while extensive, may also not adequately cover specific real-world applications like robotics. Minimal literature explores adapting a VQGAN tokenizer to domain-specific data. In this work, as our tokenization decouples dynamics information from context conditions, we hypothesize that while our model may encounter unseen objects like different robot types in downstream tasks, the fundamental knowledge of physics—such as motions and interactions—learned by the transformer from diverse scenes is commonly shared. This hypothesis is supported by our experiments transferring iVideoGPT from mixed pre-training data to the unseen BAIR dataset [20], where the pre-trained transformer can zero-shot generalize to predict natural motions, requiring only the tokenizer to be fine-tuned for unseen robot grippers (see Figure 7). This property is particularly important for scaling GPT-like transformers to large sizes, enabling lightweight alignment across domains while keeping the transformer intact. We leave an in-depth analysis of tokenizer adaptation for future work.
@@ 4 Experiments
In this section, we evaluate iVideoGPT on three different settings and compare its performance with prior state-of-the-art methods. We demonstrate that iVideoGPT is versatile to provide competitive performance across a range of tasks (Section 4.1, 4.2, and 4.3) and conduct in-depth analysis to understand the prediction ability, data efficiency, model scaling, and computational efficiency (Section 4.4). Experimental details can be found in Appendix A.
@@ Results.
Figure 4 presents the success rates of iVideoGPT compared to baseline models. While Tian et al. [82] observed that excellent perceptual metrics do not always correlate with effective control performance, iVideoGPT outperforms all baselines in two RoboDesk tasks with a large margin and achieves comparable average performance to the strongest model, SVG [87]. In Appendix B.2, we analyze iVideoGPT’s suboptimal performance on the open slide task, which is attributed to the uncertainty inherent in low-resolution observations, exacerbated by the discrete tokenization.
@@ 4.3 Visual Model-based Reinforcement Learning
We conduct experiments on six robotic manipulation tasks of varying difficulty from Meta-World [101]. Leveraging iVideoGPT as interactive world models, we have developed a model-based RL method adapted from MBPO [40], which augments the replay buffer with synthetic rollouts to train a standard actor-critic RL algorithm (see Appendix A.5 for the pseudo-code). Our implementation builds upon DrQ-v2 [99], a state-of-the-art visual model-free RL method. We also compare against a state-of-the-art model-based RL algorithm, DreamerV3 [31].
@@ Results.
Figure 6 shows that our model-based algorithm not only remarkably improves the sample efficiency over its model-free counterpart but also matches or exceeds the performance of DreamerV3. To our knowledge, this marks the first successful application of MBPO to visual continuous control tasks. These results highlight the opportunity, with powerful world models, to eliminate the need for latent imagination—a common strategy used in advanced MBRL systems to train policies on rollouts of latent states within world models [28, 77] (see comparison in Figure 5). Our development of performant MBRL algorithms decoupling model and policy learning can substantially simplify the design space, thereby greatly enhancing the practicality and effectiveness of MBRL algorithms in real-world applications.
Figure 6: Visual model-based RL on Meta-world. We report the mean and 95% confidence interval across five runs, measuring success rates over 20 evaluation episodes. PT denotes pre-training.
@@ 4.4 Model Analysis
Large-scale pre-trained models have proven effective, especially in data-scarce scenarios. Figure 8(a) shows iVideoGPT’s performance when fine-tuned with various data sizes on the BAIR dataset. We observe that pre-training offers minimal benefits when full downstream data is available, yet the advantages become significant under data scarcity (with 100 or 1,000 trajectories). The fast adaptation ability with a handful of data is particularly crucial in model-based RL. As presented in Figure 6, world models trained from scratch may generate inaccurate predictions, thereby degenerating the sample efficiency that is vital for model-based agents.
@@
Tokenization.
Finally, we evaluate the effectiveness of our compressive tokenization by comparing it against standard VQGAN tokenizers that independently convert each frame into $16\times 16$ and $4\times 4$ tokens. We train three tokenizers from scratch on RoboNet for the same number of steps. As Figure 8(c), the tokenizer with $4\times 4$ tokens suffers from low reconstruction quality due to its insufficient capacity. Our proposed tokenization method slightly compromises reconstruction quality compared to the standard $16\times 16$ tokenizer but can provide more consistent contextual information, which is beneficial for video prediction tasks. Additionally, it significantly enhances computational efficiency by 16$\times$ measured by the number of an autoregressive transformer’s forward passes.
@@
(c) Tokenization
Figure 8: Model analysis. (a) Video prediction results with various fine-tuning strategies and data sizes on BAIR. (b) Validation losses for the 138M and 436M models on the pre-training dataset. (c) Reconstruction quality of different tokenizers on RoboNet ($64\times 64$ resolution).

### v3

@@ Abstract
World models empower model-based agents to interactively explore, reason, and plan within imagined environments for real-world decision-making. However, the high demand for interactivity poses challenges in harnessing recent advancements in video generative models for developing world models at scale. This work introduces Interactive VideoGPT (iVideoGPT), a scalable autoregressive transformer framework that integrates multimodal signals—visual observations, actions, and rewards—into a sequence of tokens, facilitating an interactive experience of agents via next-token prediction. iVideoGPT features a novel compressive tokenization technique that efficiently discretizes high-dimensional visual observations. Leveraging its scalable architecture, we are able to pre-train iVideoGPT on millions of human and robotic manipulation trajectories, establishing a versatile foundation that is adaptable to serve as interactive world models for a wide range of downstream tasks. These include action-conditioned video prediction, visual planning, and model-based reinforcement learning, where iVideoGPT achieves competitive performance compared with state-of-the-art methods. Our work advances the development of interactive general world models, bridging the gap between generative video models and practical model-based reinforcement learning applications. Code and pre-trained models are available at https://thuml.github.io/iVideoGPT.
@@ 1 Introduction
Despite the fundamental connection, significant gaps remain between generative models for video generation and visual world models for agent learning. One primary challenge is achieving the best of both interactivity and scalability. In model-based RL, world models predominantly utilize recurrent network architecture. This design naturally allows the transition of observations or latent states conditioned on actions in each step, facilitating interactive behavior learning [29, 80, 34]. However, these recurrent models mostly focus on games or simulated environments with simple visuals and have limited capability to model complex, in-the-wild data at scale [48, 81]. On the other hand, Internet-scale video generative models [37, 7, 11] can synthesize realistic long videos that are controllable via text descriptions [109] or future action sequences [101] at the beginning of generation. Although suitable for high-level planning [19], their trajectory-level interactivity does not provide sufficient granularity needed by agents to intervene step-by-step during the simulation to learn precise basic skills efficiently. This dilemma naturally raises the question:
How can we leverage the advancements in scalable video generative models for developing interactive visual world models?
@@
In this work, we explore world models that are both interactive and scalable within a GPT-like autoregressive transformer framework [90, 75]. Pioneering efforts have been made recently through diffusion models [102] and masked generative models [12]. Nevertheless, utilizing autoregressive transformers offers distinct advantages such as seamless integration with the established Large Language Model (LLM) ecosystem [110] and greater flexibility in handling diverse conditions without the need for specific architectural modifications like adapter modules [77, 107]. We present Interactive VideoGPT (iVideoGPT), a scalable world model architecture that incorporates multimodal signals, including visual observations, actions, and rewards, in an interactively autoregressive manner. Unlike multimodal LLMs that discretize visual observations into tokens frame-by-frame using image tokenizers [55], a key innovation of iVideoGPT for enhancing scalability is to learn compressive tokenization for each observation conditioned on rich contextual observations, achieving an asymptotic 16$\times$ reduction in token sequence length. We highlight that more compact video tokenization could not only facilitate more efficient training and generation but also enhance video quality. This is achieved by decoupling context from dynamics, allowing the model to focus on predicting the motion of objects while maintaining temporal consistency within the scene [99].
@@
We demonstrate a series of practical applications of iVideoGPT for visual robotic manipulation, as illustrated in Figure 1. Mirroring the two-phase approach popularized by LLMs, our method involves pre-training followed by domain-specific adaptation. During pre-training, iVideoGPT is scalable for action-free video prediction across a mixture of over one million robotic and human manipulation trajectories [70, 25]. The pre-trained iVideoGPT serves as a single, adaptable foundation of interactive world models for various downstream tasks, such as action-conditioned video prediction [21, 16], visual planning [86], and visual model-based RL [105]. Additionally, we showcase the pre-trained transformer’s preliminary zero-shot video generation capability without fine-tuning, requiring only tokenizer adaptation for unseen domains. We further explore a variant of iVideoGPT for goal-conditioned video prediction, underscoring the flexibility of sequence modeling.
@@
Figure 1: Practical applications of iVideoGPT, which is designed to scale, allowing pre-training on millions of human and robotic manipulation trajectories. This results in a single, versatile foundation of interactive world models, adaptable to a wide range of downstream tasks.
@@
  • We pre-train iVideoGPT on a large-scale dataset comprising millions of robotic and human manipulation trajectories and adapt it to domain-specific tasks. The pre-trained models have been publicly available to encourage further research.
@@ 2 Problem Formulation
As depicted in Figure 2, a majority of advanced video generation models [101, 8, 104], including VideoGPT, can not deal with the interactive video prediction problem because they design non-causal modules fusing information along the temporal dimension, lacking the ability for causal, intermediate action control during generation (see extended discussion in Appendix C.2). Existing world models in the literature of MBRL [29, 80], such as Dreamer, utilize recurrent architecture but lack scalability.
@@ 3 Interactive VideoGPT
In this section, we introduce Interactive VideoGPT, a scalable world model architecture with great flexibility to integrate multimodal signals, including visual observations, actions, rewards, and other potential sensory inputs. At its core, iVideoGPT consists of a compressive tokenizer to discretize video frames and an autoregressive transformer predicting subsequent tokens (Section 3.1). This model can acquire common knowledge of motions and interactions in various scenes through pre-training on diverse human and robotic manipulation videos (Section 3.2) and then effectively transfer to downstream tasks incorporating additional modalities (Section 3.3).
@@ 3.1 Architecture
After tokenization, the video is flattened into a sequence of tokens: $x=(z_{1}^{(1)},\dots,z_{1}^{(N)},\text{{[S]}},z_{2}^{(1)},\dots,z_{2}^{(N)},\dots,\text{{[S]}},z_{T_{0}+1}^{(1)},\dots,z_{T_{0}+1}^{(n)},\dots)$ with a length of $L=(N+1)T_{0}+(n+1)(T-T_{0})-1$. Special slot tokens [S] are inserted to delineate frame boundaries and facilitate the integration of extra low-dimensional modalities such as actions (see Section 3.3 for details). As Figure 3b, a GPT-like autoregressive transformer is utilized for interactive video prediction through next-token generation frame-by-frame. In this work, we take the model size of GPT-2 [76] but adopt the LLaMA architecture [87] in order to embrace the latest innovations for LLM architecture, applying pre-normalization using RMSNorm [106], SwiGLU activation function [83], and rotary positional embeddings [85].
@@ 3.2 Pre-Training
While there are numerous videos available on the Internet, due to computational limitations, we specifically pre-train iVideoGPT for the robotic manipulation domain. We leverage a mixture of 35 datasets from the Open X-Embodiment (OXE) dataset [70] and the Something-Something v2 (SSv2) dataset [25], totaling 1.4 million trajectories (see Appendix A.2 for details). OXE is a diverse collection of robot learning datasets from a variety of robot embodiments, scenes, and tasks. These datasets are highly heterogeneous but can be easily unified in the action-free video prediction task. To further enhance the diversity, we also include SSv2, a dataset of human-object interaction videos, as previous work has demonstrated knowledge transfer from these human manipulation videos for learning a world model for robotic manipulation tasks [99, 62].
Flexibility of sequence modeling.
A sequence of tokens provides a flexible way to specify tasks, inputs, and outputs [60, 76]. To preliminarily showcase this flexibility, we introduce a variant of iVideoGPT for goal-conditioned video prediction: $p(o_{T_{0}+1:T}|o_{1:T_{0}},o_{T})$, where the model predicts a video sequence reaching a specified goal observation $o_{T}$. This is simply achieved by rearranging the frame sequence as $\tilde{o}_{1:T}=(o_{T},o_{1},o_{2},\dots,o_{T-1})$ while keeping the architecture and training procedure consistent as above (see details in Appendix A.2). Qualitative results of goal-conditioned prediction are shown in Figure 4, with further exploration left for future work [^2] Unless otherwise specified, action- and goal-free video prediction is used as the default pre-training objective to obtain pre-trained models for all experiments..
@@ 3.3 Fine-Tuning
Our architecture is also designed to flexibly incorporate additional modalities for learning interactive world models, as illustrated in Figure 3b. Actions are integrated by linear projection and adding to the slot token embeddings. For reward prediction, instead of learning independent reward predictors, we add a linear head to the last token’s hidden state of each observation. This multi-task learning approach can enhance the model’s focus on task-relevant information, thereby improving prediction accuracy for control tasks [57]. We use a mean-squared error loss for reward prediction in addition to the cross-entropy loss in Eq. (3).
@@
We choose to update the full model, including the tokenizer, for downstream tasks, finding this strategy more effective than parameter-efficient fine-tuning methods [39]. This is likely due to the limited diversity of our pre-trained data compared to Internet-scale images, which, while extensive, may also not adequately cover specific real-world applications like robotics. Minimal literature explores adapting a VQGAN tokenizer to domain-specific data. As our tokenization is designed for decoupling dynamics information from context conditions, we hypothesize that while our model may encounter unseen objects like different robot types in downstream tasks, the fundamental knowledge of physics—such as motions and interactions—learned by the transformer from diverse scenes is commonly shared. This hypothesis is supported by our experiments transferring iVideoGPT from mixed pre-training data to the unseen BAIR dataset [21], where the pre-trained transformer can zero-shot generalize to predict natural motions, requiring only the tokenizer to be fine-tuned for unseen robot grippers (see Figure 8). This property is particularly important for scaling GPT-like transformers to large sizes, enabling lightweight alignment across domains while keeping the transformer intact. We leave an in-depth analysis of tokenizer adaptation for future work.
@@ 4 Experiments
In this section, we evaluate iVideoGPT in three different control-relevant settings and compare its performance with prior state-of-the-art methods. We demonstrate that iVideoGPT is versatile to provide competitive performance across a range of tasks (Section 4.1, 4.2, and 4.3) and conduct in-depth analysis to understand the tokenization and prediction ability, data efficiency, model scaling, and computational efficiency (Section 4.4). Experimental details can be found in Appendix A.
@@ Results.
Figure 5 presents the success rates of iVideoGPT compared to baseline models. While Tian et al. [86] observed that excellent perceptual metrics do not always correlate with effective control performance, iVideoGPT outperforms all baselines in two RoboDesk tasks with a large margin and achieves comparable average performance to the strongest model, SVG [91]. In Appendix C.3, we analyze iVideoGPT’s suboptimal performance on the open slide task, which is attributed to both limitations of discretization in our model and imperfect built-in reward design of the benchmark.
@@ 4.3 Visual Model-based Reinforcement Learning
We conduct experiments on six robotic manipulation tasks of varying difficulty from Meta-World [105]. Leveraging iVideoGPT as interactive world models, we have developed a model-based RL method adapted from MBPO [42], which augments the replay buffer with synthetic rollouts to train a standard actor-critic RL algorithm (see Appendix A.5 for the pseudo-code). Our implementation builds upon DrQ-v2 [103], a state-of-the-art visual model-free RL method. We also compare against a state-of-the-art model-based RL algorithm, DreamerV3 [32], with and without world model pre-training [81].
@@ Results.
Figure 7 shows that our model-based algorithm not only remarkably improves the sample efficiency over its model-free counterpart but also matches or exceeds the performance of DreamerV3. To our knowledge, this reports the first successful application of MBPO to visual continuous control tasks. These results highlight the opportunity, with powerful world models, to eliminate the need for latent imagination—a common strategy used in advanced MBRL systems to train policies on rollouts of latent states within world models [29, 80] (see comparison in Figure 6). Our development of performant MBRL algorithms decouples model and policy learning, where iVideoGPT simply serves as a drop-in replacement of the environment. This can substantially simplify the design space, thereby greatly enhancing the practicality and effectiveness of MBRL algorithms in real-world applications.
Comparison to recurrent world models.
We argue that recurrent world models lack the capacity for large-scale pre-training on real-world data—a crucial capability for modern foundation models. To validate this, we pre-train DreamerV3 XL (200M parameters, comparable to iVideoGPT) on the same dataset. As shown in Figure 11 in the Appendix, DreamerV3 fails to capture natural robot dynamics, yielding low-quality, blurred predictions. Further evaluation on the Meta-World benchmark in Figure 7 reveals that DreamerV3 cannot benefit from such ineffective pre-training.
Figure 7: Visual model-based RL on Meta-world. (Left) Aggregated results report interquartile mean and 95% confidence interval (CI) [2] across a total of 30 runs over six tasks. (Right) Individual results for each task, report mean and 95% CI across five runs, measuring success rates over 20 evaluation episodes. PT denotes pre-training.
@@ 4.4 Model Analysis
Large-scale pre-trained models have proven effective, especially in data-scarce scenarios. Figure 9(a) shows iVideoGPT’s performance when fine-tuned with various sizes of action-free BAIR trajectories. We observe that pre-training offers minimal benefits when full downstream data is available, yet the advantages become significant under data scarcity (with 100 or 1,000 trajectories). We also adapt iVideoGPT using 1,000 action-conditioned BAIR trajectories, achieving an FVD of 82.3. The fast adaptation ability with a handful of data is particularly crucial in model-based RL. As shown in Figure 7, world models trained from scratch may generate inaccurate predictions, thereby degenerating the sample efficiency that is vital for model-based agents.
@@
Tokenization efficiency.
We evaluate the effectiveness of our compressive tokenization by comparing it against standard VQGAN tokenizers that independently convert each frame into $16\times 16$ and $4\times 4$ tokens. We train three tokenizers from scratch on RoboNet for the same number of steps. As Figure 9(c), the tokenizer with $4\times 4$ tokens suffers from low reconstruction quality due to its insufficient capacity. Our proposed tokenization method slightly compromises reconstruction quality compared to the standard $16\times 16$ tokenizer but can provide more consistent contextual information, which is beneficial for video prediction tasks. More importantly, it significantly enhances computational efficiency with a significantly fewer amount of tokens, which greatly saves time and memory, allowing us to scale the model size with fewer costs (see quantitative results in Appendix B.5).
@@
(c) Tokenization efficiency
Figure 9: Model analysis. (a) Video prediction results with various fine-tuning strategies and data sizes on BAIR. (b) Validation losses for the 138M and 436M transformer models on the pre-training dataset. (c) Computational efficiency and reconstruction quality of different tokenizers.
Context-dynamics decoupling.
Our tokenizer is designed with a bottleneck of much fewer tokens, focusing only on capturing necessary dynamics information for future frames while sharing contextual information with initial frames to reconstruct raw pixels. To explicitly visualize this decoupling of context and dynamics information, we drop cross-attention blocks to context frames in the decoder when reconstructing future frames. The results in Figure 10 show that the decoder can still reproduce the movement trajectories accurately but with minimal contextual information. This visualization supports the explanation of our model’s generalization capability shown in Figure 8.
Goal-conditioned prediction.
In Figure 4, we also showcase video prediction generated by goal-conditioned iVideoGPT, pre-trained on massive human and robotic videos (Section 3.2). Unlike action-free prediction, which often results in trajectories diverging from the ground truth, the goal-conditioned model produces more accurate paths to reach specified goals. We believe this highlights the potential of leveraging the flexibility of a unified sequence modeling paradigm.
Figure 10: Context-dynamics decoupling in our compressive tokenization. By removing cross-attention from future frames to context frames, the decoder can still reconstruct a trajectory that moves in the same way as the original, but the visual context is almost entirely missing.